OpenStudy (anonymous):

the slope of the line perpendicular to another line with the equation 2y+3x-4=0 is -3/2 . true of false?

OpenStudy (saifoo.khan):

False.

OpenStudy (anonymous):

your equation could be presented as: 2y=-3x+4 y=-3/2 x +2 slope of this line is m=-3/2 so, the line perpendicular to this one should be: 2/3 => statement is false